OpenStudy (anonymous):

Solve. 7x – 5 = 74x + 7 @mathematics

OpenStudy (akashdeepdeb):

These are simultaneous equations Solve them by taking the variables to one side and the constants to the other. 7x – 5 = 74x + 7 -12 = 67x x = -12/67 Got it? :)

OpenStudy (anonymous):

@AkashdeepDeb im sorry its Solve. 7x – 5 = 7^4x + 7

OpenStudy (anonymous):

IS IT 7x – 5 = (7^4)x + 7 OR THE WAY U GAVE?